Gayle is thinking of starting her own business. Total fixed costs are $19,000 per month and unit variable costs are estimated at $37.50. From some preliminary studies that she completed, she forecasts sales of 1,400 units at $50 each, 1,850 units at $48 each, 2,500 units at $46 each, and 2,750 units at $44 each. What price would you recommend Gayle set for her products? 


1
Expert's answer
2021-12-15T06:15:10-0500

When price = "\$50" then sales =1400 units


Revenue = (sales) "\u00d7" (price) (1400) "\u00d7" (50) = "\$70,000"


cost = (variable) + (fixed)

=1400 (37.5) + 19,000 = "\$71,500"


Profit= (Revenue) - (Cost)

= "-\$70000-\$71,500"

="-\$1500"


Hence,Loss = "\$1500"


when price = "\$48" sales = 1850 Revenue = (1850) "\u00d7" (48) ="\$88,800"

Cost = 1850 (37-5) + 19,000 ="\$88,375"

Profit= "\$88,800 - \$88,375 = \$425"

Profit = "\$425"


When price = "\$46" , sales = 2500

Revenue = (2500)x (46) ="\$115,000"

cost = 2500(37.5) + 19,000 ="\$112,750"

Profit;"\$115,000 - \$112,750 =\$2,250"

Profit= "\$2,250"


When price = "\$44" , sales = 2750

Revenue = (2750)x (44) ="\$121,000"

cost = 2750(37.5) + 19,000 ="\$122,125"

Profit="\$121,000-\$122,125 =-\$1,125" Loss = "\$1,125"


He should choose a price of "\$46" since it gives the most profit